Find the charge required to flow through an electrolyte to liberate on atom of (a) a monovalent material and (b) a divalent material.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. (a) 1eq. mass of the substance requires 96500 coulombs Since the element is monoatomic, thus eq. mass = mol. Mass 6.023 × 10^23 atoms require 96500 C 1 atoms require 96500/6.023 *10^23 C = 1.6 * 10^-19 C (b) Since the element is diatomic eq.mass = (1/2) mol.mass ∴ (1/2) * 6.023 * 10^23 atoms 2eq. 96500C ⇒ 1 atom require = 96500 * 2/6.023 * 10^23 = 3.2 * 10^-19 C